# Two-generation of traceless matrices over finite fields

### v2

@@ 1. Introduction
This $2$-generation phenomenon is well-established across various classical algebraic structures over classical fields such as the complex numbers. For simple associative algebras, this property is underscored by the classical Burnside irreducibility theorem, which posits that over algebraically closed fields, any two matrices without a common proper invariant subspace generate the whole matrix algebra. Similarly, for simple algebraic groups, it is a result of Guralnick [Gur98, Theorem 3.3] that over an algebraically closed field of any characteristic, a simple algebraic group $G$ is topologically $3/2$-generated, meaning that for every nontrivial $x$ there exists $y$ so that $\langle x,y\rangle$ is dense in $G$. Furthermore, simple Lie algebras over any field of characteristic 00 have been shown to be $2$-generated by Kuranishi [Kur51] (see also [BO76] for a more generic approach), and explicit generating pairs were found by Grozman and Leites [LG96]. This was extended to $3/2$-generation in characteristic 00 by Ionescu [Ion76] and in positive characteristic $p>3$ by Bois [Boi09]. More recently, Chistopolskaya [Chi18] has explored nilpotent pairs of generators in traceless matrices.
@@ 3.1. Bounded characteristic and large field
3.2. Large characteristic
Whenever we have a consistent set $\{\lambda_{1},\lambda_{2},\dots,\lambda_{n}\}$ with entries in the integers, we can project it modulo $p$. The projected set is consistent as long as $p>4\max\{|\lambda_{1}|,|\lambda_{2}|,\dots,|\lambda_{n}|\}$, in which case the associated diagonal matrix $\operatorname{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ together with $\mathbf{1}$ modulo $p$ generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{p})$ over $\mathbf{F}_{p}$ and therefore also generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ over $\mathbf{F}_{q}$.
For example, the consistent matrix obtained from the set (2) gives a generating pair for $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ as long as $p>2^{n+1}$, a bound that grows exponentially in $n$. We can do better.
Consistent sets with entries in positive integers are intimately related to the concept of distinct sum sets [GS80]. These are sets of integers $\{a_{i}\}_{i=0}^{n}$ satisfying
$0=a_{0}<a_{1}<a_{2}<\cdots<a_{n}$
with the property that the sums $a_{i}+a_{j}$ for $i<j$ represent each integer at most once. Given such a set, consider
$S=\{a_{1},a_{2},\dots,a_{n},-\sum_{i=1}^{n}a_{i}\}.$
Note that all of the $n+1$ elements of $S$ are distinct and their sum is 00. Since $a_{j}-\sum_{i=1}^{n}a_{i}<0$ for any $j$, the set $S$ is consistent. Therefore a distinct sum set gives a consistent set $S$ of the same size. Moreover, all elements of $S$ are of absolute value at most $n\cdot a_{n}$. On the other hand, given any set of integers that is consistent modulo $p$, one gets a distinct sum set in $\mathbf{F}_{p}$ by adding the element 00 to it.
Much is known about existence of distinct sum sets. It is shown in [GS80, Theorem 1 (3)] that given $n$, it is possible to construct a distinct sum set whose every element is of size at most $n^{2}+O(n^{36/23})$. The associated consistent set will therefore consist of elements whose absolute values are $O(n^{3})$, giving in turn consistent matrices modulo $p$ for a polynomially large $p$ in terms of $n$.
Proposition 3.3.
Let $p=\Omega(n^{3})$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@ 4.1. Polynomials with consistent roots
Let $\mathbf{F}$ be a splitting field for $f$. Let $C$ be the companion matrix of $f$. Since the roots of $f$ form a consistent set, we have $C\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ and the eigenvalues of $C$, being the roots of $f$, are all distinct in $\mathbf{F}$. Therefore $C$ diagonalizes to a consistent matrix $D=PCP^{-1}$ for some $P\in\operatorname{GL}_{n}(\mathbf{F})$, which generates $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ together with the matrix $\mathbf{1}$. Thus, according to [Boi09, Proposition 1.1.3], there exist $n^{2}-1$ Lie monomials $\{g_{i}(X,Y)\}_{i}$ in two variables whose evaluations in $(D,\mathbf{1})$ span the vector space $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Since Lie polynomials are equivariant, [^1] Meaning that if $g(X,Y)$ is a Lie polynomial, then $P^{-1}g(X,Y)P=g(P^{-1}XP,P^{-1}YP)$. the matrices $g_{i}(C,P^{-1}\mathbf{1}P)=P^{-1}g_{i}(D,\mathbf{1})P$ also span $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Now, both $C$ and $P^{-1}\mathbf{1}P$ belong to $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and as each $g_{i}$ is a Lie monomial, its evaluation $g_{i}(C,P^{-1}\mathbf{1}P)$ is in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. Therefore $C$ and $P^{-1}\mathbf{1}P$ generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ as a Lie algebra over $\mathbf{F}_{q}$. ∎
@@ 5. Even characteristic
For instance, if $\alpha\in\mathbf{F}_{q^{n}}$ is a sharply traceless element, $\operatorname{diag}(\alpha,\operatorname{Fr}(\alpha),\dots,\operatorname{Fr}^{n-1}(\alpha))$ is semiconsistent in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q^{n}})$ as long as $n\neq 4$.
@@ 6. Exceptional cases
For any $x,y\in\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$, we have
@@
in $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$. We multilinearize this identity by considering it in the Lie algebra $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})\otimes_{\mathbf{F}}\mathbf{F}[T]/(T^{2})$ and using it with the element $Tx+y$ for any $x,y\in\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$. [^6] We thank Alexander Premet who showed us this trick. It substantially shortens our previous argument. Note that, by Jacobson’s formula [Jac79],
@@
The case $(n,p)=(4,2)$ is proved with the same method.
@@
There are forms $a,b,c,d$ on $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ so that for any $x,y\in\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$, we have
@@
Let $x\in\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$. The characteristic polynomial of $x$ is $t^{4}+\alpha(x)t^{2}-\beta(x)t+\det(x)$, where $\alpha,\beta$ are forms in $x$ [^7] These forms are computable as elementary symmetric polynomials $e_{2},e_{3}$ in the eigenvalues of $x$. As we are in characteristic is $2$, these polynomials are not expressible in terms of power sums, so the formula is not as nice as the one in the previous section. Therefore we have
@@
in $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$. As in odd characteristic, multilinearize this identity by using it with the element $Tx+y$ with $T^{2}=0$. By Jacobson’s formula, we have
@@
Note that $\alpha(Tx+y),\beta(Tx+y)$ are polynomials in $T$, since they are expressible as coefficients of the polynomial $\det(tI-Tx-y)$ in $t$. Write $\alpha(Tx+y)=c+Td$ in $\mathbf{F}[T]/(T^{2})$ and similary $\beta(Tx+y)=a+Tb$, where $a,b,c,d$ are forms on $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ evaluated at $x,y$. By inserting $Tx+y$ into (7), we thus obtain

### v3

@@ 1. Introduction
This $2$-generation phenomenon is well-established across various classical algebraic structures over classical fields such as the complex numbers. For simple associative algebras, this property is underscored by the classical Burnside irreducibility theorem, which posits that over algebraically closed fields, any two matrices without a common proper invariant subspace generate the whole matrix algebra. Similarly, for simple algebraic groups, it is a result of Guralnick [Gur98, Theorem 3.3] that over an algebraically closed field of any characteristic, a simple algebraic group $G$ is topologically $3/2$-generated, meaning that for every nontrivial $x$ there exists $y$ so that $\langle x,y\rangle$ is dense in $G$. Furthermore, simple Lie algebras over any field of characteristic 00 have been shown to be $2$-generated by Kuranishi [Kur51] (see also [BO76] for a more generic approach), and explicit generating pairs were found by Grozman and Leites [LG96]. This was extended to $3/2$-generation in characteristic 00 by Ionescu [Ion76] and in infinite fields of positive characteristic $p>3$ by Bois [Boi09]. More recently, Chistopolskaya [Chi18] has explored nilpotent pairs of generators in traceless matrices.
@@ 3. Bounded rank
[…]
Consistent sets with entries in positive integers are intimately related to the concept of distinct sum sets [GS80]. These are sets of integers $\{a_{i}\}_{i=0}^{n}$ with
$0=a_{0}<a_{1}<\dots<a_{n}$
and the property that the sums $a_{i}+a_{j}$ for $i<j$ represent each integer at most once. From such a set, one obtains the consistent set of integers
$S=\{a_{1},a_{2},\dots,a_{n},-\sum_{i=1}^{n}a_{i}\},$
all of whose elements are of absolute value at most $n\cdot a_{n}$.
Much is known about existence of distinct sum sets. It is shown in [GS80, Theorem 1 (3)] that given $n$, one can construct such a set whose every element is of size at most $n^{2}+O(n^{36/23})$, and therefore the associated consistent set consists of elements with absolute values bounded by $O(n^{3})$. As a consequence, for $p=\Omega(n^{3})$, the Lie algebra $\mathfrak{sl}_{n}(\mathbb{F}_{q})$ is $2$-generated.
@@ 4.1. Polynomials with consistent roots
Let $\mathbf{F}$ be a splitting field for $f$. Let $C$ be the companion matrix of $f$. Since the roots of $f$ form a consistent set, we have $C\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ and the eigenvalues of $C$, being the roots of $f$, are all distinct in $\mathbf{F}$. Therefore $C$ diagonalizes to a consistent matrix $D=PCP^{-1}$ for some $P\in\operatorname{GL}_{n}(\mathbf{F})$, which generates $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ together with the matrix $\mathbf{1}$. Thus, according to [Boi09, Proposition 1.1.3], there exist $n^{2}-1$ Lie monomials $\{g_{i}(X,Y)\}_{i}$ in two variables whose evaluations in $(D,\mathbf{1})$ span the vector space $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Since Lie polynomials are equivariant (meaning that if $g(X,Y)$ is a Lie polynomial, then $P^{-1}g(X,Y)P=g(P^{-1}XP,P^{-1}YP)$) the matrices $g_{i}(C,P^{-1}\mathbf{1}P)=P^{-1}g_{i}(D,\mathbf{1})P$ also span $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Now, both $C$ and $P^{-1}\mathbf{1}P$ belong to $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and as each $g_{i}$ is a Lie monomial, its evaluation $g_{i}(C,P^{-1}\mathbf{1}P)$ is in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. Therefore $C$ and $P^{-1}\mathbf{1}P$ generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ as a Lie algebra over $\mathbf{F}_{q}$. ∎
@@ 5. Even characteristic
For instance, if $\alpha\in\mathbf{F}_{q^{n}}$ is a sharply traceless element, the diagonal matrix $\operatorname{diag}(\alpha,\operatorname{Fr}(\alpha),\dots,\operatorname{Fr}^{n-1}(\alpha))$ is semiconsistent in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q^{n}})$ as long as $n\neq 4$.
@@ 6.1. $(n,p)=(3,3)$
For any $x,y\in\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$, the following identity holds in $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$:
@@
in $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$. This identity holds over any field of characteristic $3$, as well as any subquotient of such a field. We can thus multilinearize this identity by using it in the Lie algebra $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})\otimes_{\mathbf{F}}\mathbf{F}[T]/(T^{2})$ with the element $Tx+y$ for any $x,y\in\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$. [^5] We thank Alexander Premet who showed us this trick. It substantially shortens our previous argument. Note that, by Jacobson’s formula [Jac79],
@@
In fact, when the $u$-invariant [^6] The largest integer $u(\mathbf{F})\geq 0$ such that there exists an anisotropic quadratic form of dimension $u(\mathbf{F})$ over $\mathbf{F}$, or is infinity if no such largest integer exists. of $\mathbf{F}$ is at most $2$, the identity in Proposition 6.2 further implies that for generic choices of $x$ and $y$, the Lie algebra they generate is isomorphic to $\operatorname{\mathfrak{sl}}_{2}(\mathbf{F})$. This occurs, for example, when $\mathbf{F}$ is finite or algebraically closed (see [EKM08, Example 36.2]).
Corollary 6.3.
[…]
Proof.
Suppose first that $G$ is nondegenerate. Hence $x,y,[x,y]$ are linearly independent. The restriction of the trace form to the span of $x,y,[x,y]$ is thus a nondegenerate quadratic form of dimension $3$. This form is isotropic, so by the Witt decomposition theorem, it contains a hyperbolic plane spanned by some $e,f$. Thus $\operatorname{Tr}(e^{2})=\operatorname{Tr}(f^{2})=0$ and $\operatorname{Tr}(ef)=1$. Let $h=[e,f]$. Using Proposition 6.2, we then compute $[h,e]=2e$ and $[h,f]=-2f$ in $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$. These formulas further imply that $e,f,h$ are linearly independent. After passing to $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$, the Lie algebra generated by $x$ and $y$ is thus isomorphic to $\operatorname{\mathfrak{sl}}_{2}(\mathbf{F})$.
Assume now that $G$ is degenerate. We can assume that the Lie algebra generated by the images of $x,y$ in $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$ is of dimension $3$. Hence $x,y,[x,y]$ are linearly independent and the restriction of the trace form to their span $L=\langle x,y,[x,y]\rangle$ is degenerate. Thus there is a nonzero $z\in L$ with $\operatorname{Tr}(zx)=\operatorname{Tr}(zy)=\operatorname{Tr}(z[x,y])=0$. Using Proposition 6.2, we conclude that $\operatorname{ad}_{z}$ is nilpotent of degree $2$ on $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$. The Lie algebra $\operatorname{\mathfrak{sl}}_{2}(\mathbf{F})$, however, does not contain nonzero nilpotent elements of degree $2$ when $\mathbf{F}$ is of odd characteristic. This completes the proof. ∎
[…]
6.2. $(n,p)=(4,2)$
This case is handled using the same method.
@@ 6.2. $(n,p)=(4,2)$
There are forms $a,b,c,d$ on $\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$ so that for any $x,y\in\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$, the following identity holds in $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$:
@@
Let $x\in\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$. The characteristic polynomial of $x$ is $t^{4}+\alpha(x)t^{2}-\beta(x)t+\det(x)$, where $\alpha,\beta$ are forms in $x$ with coefficients in $\mathbf{F}_{2}$. [^8] These forms are computable as elementary symmetric polynomials $e_{2},e_{3}$ in the eigenvalues of $x$. As we are in characteristic $2$, these polynomials are not expressible in terms of power sums, so the formula is not as clear as the one in the previous section. Therefore we have
@@
in $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$. As in the proof of Proposition 6.2, we multilinearize this identity by using it with the element $Tx+y$ with $T^{2}=0$. By Jacobson’s formula, we have
@@
Note that $\alpha(Tx+y),\beta(Tx+y)$ are polynomials in $T$, since they are expressible as coefficients of the polynomial $\det(tI-Tx-y)$ in $t$. Write $\alpha(Tx+y)=c+Td$ in $\mathbf{F}[T]/(T^{2})$ and similarly $\beta(Tx+y)=a+Tb$, where $a,b,c,d$ are forms on $\operatorname{\mathfrak{sl}}_{4}(\mathbf{F})$ evaluated at $x,y$. By inserting $Tx+y$ into (7), we thus obtain
@@
A similar phenomenon to Corollary 6.3 appears to hold in this case as well, although we have not been able to prove it. Generically, the Lie algebra generated by two elements in $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ seems to be isomorphic to $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$.